If the area of a rectangle is 20 cm. and its length is 19 less than 6 times its width, what are the length and width? Can you show me how you got the answer?
Answer by richard1234(7193)

Length = 5
Width = 4 (to check, 5 = 6*4 - 19)

You can guess and check at the answer in some cases. A more algebraic solution is to let W be the width, then the length is 6W-19, and W(6W-19) = 20. This is equivalent to the quadratic 6W^2 - 19W - 20 = 0. Solve via factoring or quadratic formula.